A Bibliography of Papers in Lecture Notes in Computer Science (2014): Volumes 8349–??

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.02

Title word cross-reference

4 [9].

-Round [9].

1 [21].

Achieving [7]. Adaptively [16]. against [19]. Amplification [18].

Based [10]. Basing [12]. Be [10]. Bit [19]. Bit-Wise [19]. Black [1, 15]. Black-Box [1, 15]. Box [1, 15]. Broadcast [18].

Can [10]. Characterizing [13]. Checkable [6]. Chosen [5]. Ciphertext [5]. Circuits [1]. Codes [20]. Coding [19]. Coin [10, 12]. Complete [13]. Complexity [14]. Composable [15]. Computation [13, 16, 15, 11].

Concurrent [8]. Constant [15, 7]. Constant-Round [15]. Construction [15]. Continuous [20]. Cryptographic [14]. Encoding [1]. Encryption [11]. Evasive [2]. Eve [21]. Extractability [3]. Fair [10]. Fairness [13]. Functions [2, 14, 10]. Generic [1]. Graded [1]. Hashing [21]. Impossibility [12]. Indistinguishability [4]. Key [11]. Knowledge [9, 6, 8, 7]. Leakage [7]. Leakage-Resilient [7]. linear [17]. Malicious [21]. malleable [19, 20, 8]. MPC [4]. Multi [17, 15]. Multi-linear [17]. Multi-Party [15]. Non [19, 20, 8]. Non-malleable [19, 20, 8]. Obfuscation [2, 3, 1, 4, 5]. One [10, 16, 12]. One-Sided [16]. One-Way [10, 12]. Optimally [10]. Optimally-Fair [10]. Party [13, 16, 15]. Permutations [12]. Point [5]. Power [11]. Probabilistically [6]. Proofs [6]. Protocol [15]. Proximity [6]. Public [11, 12]. Public-Coin [12]. Public-Key [11]. Resettably [9]. Resettably-Sound [9]. Resilient [7]. Round [9, 4, 15, 7]. Schemes [17]. Secret [17]. Secret-Sharing [17]. Secure [13, 4, 16, 11]. Security [5]. SHA [21]. SHA-1 [21]. Sharing [17]. Sided [16]. Sound [9]. Split [19]. Split-State [19]. State [19]. Statistical [8]. Tampering [19]. Tossing [10]. Trapdoor [12]. Two [13, 4, 16]. Two-Party [13, 16]. Two-Round [4]. Variant [21]. via [1, 5]. Virtual [1]. Way [10, 12]. Wise [19]. Worst [14]. Zero [9, 6, 8, 7]. Zero-Knowledge [6, 7].

References

Brakerski:2014:VBB

[1] Zvika Brakerski and Guy N. Rothblum. Virtual black-box obfuscation for all circuits via generic graded encoding. Lecture Notes in Computer Science, 8349:1–25, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-54242-8_1/;   http://link.springer.com/content/pdf/bfm:
978-3-642-54242-8/1.pdf.

Barak:2014:OEF

[2] Boaz Barak, Nir Bitansky, Ran Canetti, Yael Tauman Kalai, Omer Paneth, and Amit Sahai. Obfuscation for evasive functions. Lecture Notes in Computer Science, 8349:26–51, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-54242-8_2/.

Boyle:2014:EO

[3] Elette Boyle, Kai-Min Chung, and Rafael Pass. On extractability obfuscation. Lecture Notes in Computer Science, 8349:52–73, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-54242-8_3/.

Garg:2014:TRS

[4] Sanjam Garg, Craig Gentry, Shai Halevi, and Mariana Raykova. Tworound secure MPC from indistinguishability obfuscation. Lecture Notes in Computer Science, 8349:74–94, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-54242-8_4/.

Matsuda:2014:CCS

[5] Takahiro Matsuda and Goichiro Hanaoka. Chosen ciphertext security via point obfuscation. Lecture Notes in Computer Science, 8349:95–120, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-54242-8_5/.

Ishai:2014:PCP

[6] Yuval Ishai and Mor Weiss. Probabilistically checkable proofs of proximity with zero-knowledge. Lecture Notes in Computer Science, 8349:121–145, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-642-54242-
8_6/.

Pandey:2014:ACR

[7] Omkant Pandey. Achieving constant round leakage-resilient zeroknowledge. Lecture Notes in Computer Science, 8349:146–166, 2014. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-54242-8_7/.

Orlandi:2014:SCN

[8] Claudio Orlandi, Rafail Ostrovsky, Vanishree Rao, Amit Sahai, and Ivan Visconti. Statistical concurrent non-malleable zero knowledge. Lecture Notes in Computer Science, 8349:167–191, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-54242-8_8/.

Chung:2014:RRS

[9] Kai-Min Chung, Rafail Ostrovsky, Rafael Pass, and Muthuramakrishnan Venkitasubramaniam . 4-round resettably-sound zero knowledge. Lecture Notes in Computer Science, 8349:192–216, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-54242-8_9/.

Dachman-Soled:2014:COF

[10] Dana Dachman-Soled, Mohammad Mahmoody, and Tal Malkin. Can optimally-fair coin tossing be based on one-way functions? Lecture Notes in Computer Science, 8349:217–239, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-54242-8_10/.

Mahmoody:2014:PPK

[11] Mohammad Mahmoody, Hemanta K. Maji, and Manoj Prabhakaran. On the power of public-key encryption in secure computation. Lecture Notes in Computer Science, 8349:240–264, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-54242-8_11/.

Matsuda:2014:IBP

[12] Takahiro Matsuda. On the impossibility of basing public-coin one-way permutations on trapdoor permutations. Lecture Notes in Computer Science, 8349:265–290, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-54242-8_12/.

Asharov:2014:TCC

[13] Gilad Asharov. Towards characterizing complete fairness in secure twoparty computation. Lecture Notes in Computer Science, 8349:291–316, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-642-54242-
8_13/.

Beimel:2014:CCW

[14] Amos Beimel, Yuval Ishai, Ranjit Kumaresan, and Eyal Kushilevitz. On the cryptographic complexity of the worst functions. Lecture Notes in Computer Science, 8349:317–342, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-54242-8_14/.

Kiyoshima:2014:CRB

[15] Susumu Kiyoshima, Yoshifumi Manabe, and Tatsuaki Okamoto. Constantround black-box construction of composable multi-party computation protocol. Lecture Notes in Computer Science, 8349:343–367, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-54242-8_15/.

Hazay:2014:OSA

[16] Carmit Hazay and Arpita Patra. One-sided adaptively secure two-party computation. Lecture Notes in Computer Science, 8349:368–393, 2014. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-54242-8_16/.

Beimel:2014:MLS

[17] Amos Beimel, Aner Ben-Efraim, Carles Padró, and Ilya Tyomkin. Multilinear secret-sharing schemes. Lecture Notes in Computer Science, 8349: 394–418, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-54242-8_17/.

Hirt:2014:BA

[18] Martin Hirt, Ueli Maurer, and Pavel Raykov. Broadcast amplification. Lecture Notes in Computer Science, 8349:419–439, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-54242-8_18/.

Cheraghchi:2014:NMC

[19] Mahdi Cheraghchi and Venkatesan Guruswami. Non-malleable coding against bit-wise and split-state tampering. Lecture Notes in Computer Science, 8349:440–464, 2014. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-54242-8_19/.

Faust:2014:CNM

[20] Sebastian Faust, Pratyay Mukherjee, Jesper Buus Nielsen, and Daniele Venturi. Continuous non-malleable codes. Lecture Notes in Computer Science, 8349:465–488, 2014. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-54242-8_20/.

Albertini:2014:MHE

[21] Ange Albertini, Jean-Philippe Aumasson, Maria Eichlseder, Florian Mendel, and Martin Schläffer. Malicious hashing: Eve’s variant of SHA-1. Lecture Notes in Computer Science, 8781:1–19, November 29, 2014. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://eprint.iacr.org/2014/694;   http://link.springer.
com/chapter/10.1007/978-3-319-13051-4_1;   https://malicioussha1.
github.io/.